OpenStudy (anonymous):

Hi sir, the functions f and g are defined for real values of x by, f:x=ln(2x-1),x>1/2 g:x=2x,x>1 express each of the following in terms of f and g i) x=x/2 ii) x=ln(4x-1) iii) x=ln(4x-1)^2. please reply as soon as possible thank u

OpenStudy (anonymous):

what the monkey does \(x=\frac{x}{2}\) mean in this context? if \(x=\frac{x}{2}\) then \(x=0\)

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

is that an equal sign?

OpenStudy (anonymous):

x->x/2 its like this

OpenStudy (anonymous):

oooh \(x\to \frac{x}{2}\)

OpenStudy (anonymous):

mm mm

OpenStudy (anonymous):

then since \(g(x)=2x\) we have \[x\to \frac{x}{2}\] is \[x\to\frac{g(x)}{4}\]

OpenStudy (anonymous):

is this a 1st solution for g(x)

OpenStudy (anonymous):

i dont know the basics of these terms, am trying 1st time

OpenStudy (anonymous):

hello satellite r u there, cant get ur transmission for signals... help me

OpenStudy (anonymous):

yeah that is the first one since \(g(x)=2x\) then \(\frac{g(x)}{4}=\frac{2x}{4}=\frac{x}{2}\)

OpenStudy (anonymous):

ok ok... 2nd n 3rd please